# Excited states of coherent harmonic qubits with long-range photon coupling and dissipation

### v1

@@ III Conclusions
The results of the previous section show that when the system is in the coherent ground state, the excitation to states of higher energy involves a phenomenon of constructive interference of the $N$ qubits. The largest transition probability is obtained when the qubits change their state in a coordinated way. It may be expected that this effect of constructive interference is more pronounced when $N$ is increased. However, the values of $N$ tested so far are quite small.
@@
Actually, the model expressed by the Hamiltonian 1, whose mathematical form is quite general, has been previously applied to a system of identical physical charges oscillating in a periodic 3D lattice. A system of this kind can condense into a coherent ground state, thanks to the growth of a trapped photonic mode of the same frequency, only if the spatial density of the oscillating charges is sufficiently large. This is necessary in order that the (negative) interaction term compensates for the positive energy of the photons. The present numerical simulations at small values of $N$ thus make sense, physically, if one assumes that the system is confined in a small volume, with periodic boundary conditions for the field.
@@ IV Codes
[…]

### v2

@@ III Discussion and Conclusions
The results of the previous sections show that when the system is in the coherent ground state, the excitation to states of higher energy involves a phenomenon of constructive interference of the $N$ qubits. The largest transition probability is obtained when the qubits change their state in a coordinated way. It may be expected that this effect of constructive interference is more pronounced when $N$ is increased. However, the values of $N$ tested so far are quite small.
@@
The model expressed by the Hamiltonian (2), whose mathematical form is quite general, has been previously applied to a system of identical physical charges oscillating in a periodic 3D lattice. A system of this kind can condense into a coherent ground state, thanks to the growth of a trapped photonic mode of the same frequency, only if the spatial density of the oscillating charges is sufficiently large. This is necessary in order that the (negative) interaction term compensates for the positive energy of the photons. The present numerical simulations at small values of $N$ thus make sense, physically, if one assumes that the system is confined in a small volume, with periodic boundary conditions for the field.
@@ IV Codes
In the numerical simulations of oscillators in [10] the dimension $n_{max}$ of the finite Hilbert space of states employed in the matrix representation of the operators is determined via a “saturation” criterium. For ex., in [10] it was found that a dimension $n_{max}\geq 35$ was approximately necessary in order to obtain values of the ground state energy independent from $n_{max}$. The total space is a tensor product with dimension $n_{max}^{N}\cdot n_{photon}$, with $n_{photon}\simeq Nn_{max}$. This means that by simulating 4, 5 or more oscillators one quickly reaches the largest dimension that can be handled by standard numerical eigenvalues methods. That is a problem, especially if one wants to obtain information about the excited states of the system.
@@
plt.ylabel("Minimum of occupation $\langle$ $\dagger Q \\$rangle\$")